# Worst-Case to Expander-Case Reductions: Derandomized and Generalized

### v1

@@ Abstract
A recent paper by Abboud and Wallheimer [ITCS 2023] presents self-reductions for various fundamental graph problems, that transform worst-case instances to expanders, thus proving that the complexity remains unchanged if the input is assumed to be an expander. An interesting corollary of their self-reductions is that, if some problem admit such reduction, then the popular algorithmic paradigm based on expander-decompositions is useless against it. In this paper, we improve their core gadget, which augments a graph to make it an expander while retaining its important structure. Our new core construction has the benefit of being simple to analyze and generalize, while obtaining the following results:
@@
  • An extension of the results to other models of computation, such as the Fully Dynamic model and the Congested Clique model. In the former, we either improve or provide an alternative approach to some recent hardness results for dynamic expander graphs, by Henzinger, Paz, and Sricharan [ESA 2022].
@@ 1 Introduction
It is possible to cook up problems for which the answer to Question 1 is negative. For instance, we can solve connectivity in constant time if the input is promised to be an expander, but it requires linear time in the worst case. Less obvious examples with such gaps are counting spanning trees [32] and computing short paths [8] [^2] Using the stronger notion of spectral expanders, which we do not consider here.. However, for many (perhaps most) interesting graph problems, the answer seems to be positive: expanders do not make the problem any easier. In other words, the expander-case is also worst-case. But how do we prove that? Let us discuss three methods and their drawbacks.
@@
    The main drawback with this approach is that we are interested in answering Question 1 even when (or rather, especially when) we have not already resolved the worst-case time complexity of our problem, in which case we do not even have a matching lower bound (e.g., the Maximum Matching problem). Another drawback is that when asking for the fine-grained complexity of problems, the existing lower bounds are usually conditioned on strong assumptions, and one may hope to get an unconditional answer to Question 1.
@@
    The second approach uses the expander decomposition method. This is a popular paradigm in recent years that suggests we can solve graph problems by (1) decomposing the graph into vertex-disjoint expanders with a small number of edges between them, (2) solving the problem on each expander separately, and (3) combining all the answers efficiently. Step (3) requires problem-specific techniques. If we can solve steps (1) and (3) for a problem, then we have effectively shown that any improvement on expanders will yield an improvement on worst-case graphs, giving a positive answer to Question 1. Efficient algorithms for computing such expander decompositions (for step (1)) are known both in the Word-RAM  [5, 38, 28, 41, 34] and in other models of computation including dynamic [38], distributed [13], and recently even in streaming [17]. Applications of this paradigm have led to many breakthroughs in recent years to problems such as Maximum Flow [14], Dynamic Connectivity [23], Gomory-Hu Trees [2], Minimum Spanning Trees [33], and Triangle Enumeration [13].
@@
    • The blowup in the number of vertices and edges in $G_{exp}$ is bounded. Namely, $|V_{exp}|\leq K$ and $|E_{exp}|\leq M$ for some $K:=K(n,m)$ and $M:=M(n,m)$.
@@
    The main contribution of Abboud and Wallheimer was to show that some fundamental problems, such as $k$-Clique Detection and Maximum Matching, admit simple Direct-WTERs. In particular, their Direct-WTERs make a graph an expander by employing a core gadget that augments it with $O(n)$ vertices and $\tilde{O}(n+m)$ random edges and then applies additional gadgets that control the solution. In particular, they obtain a near-linear blowup. Their results are surprising because such Direct-WTERs do not employ any of the heavy machinery that usually comes with expander decompositions, yet they output quantitatively better expanders: the outputs of Direct-WTERs are $\Omega(1)$-expanders (by definition), whereas expander decompositions can only produce $O(1/\log n)$-expanders (that are not as expanding) [39, 6].
@@
Motivated by the appeal of the method of Direct-WTERs towards answering Question 1 and Question 2, our goal is to develop this theory further. Toward that, we address the two main limitations that were highlighted in [3]: (1) the Direct-WTERs are randomized whereas ED-WTERs are deterministic [16], and (2) the results are restricted to the Word-RAM model, whereas expander decompositions are popular tools in other models as well. In addition, we continue their line of work by providing Direct-WTERs to additional problems. Let us motivate these two topics before stating our results formally.
@@ 1.1 Deterministic Direct-WTERs
The randomized Direct-WTERs of Abboud and Wallheimer [3] prove that $\Omega(1)$-expanders are worst-case instances of many problems if we allow algorithms to be randomized. They leave us wondering if perhaps $\Omega(1)$-expanders are truly easier for deterministic algorithms. We remark that while it is believed that all algorithms can be derandomized by incurring a small polynomial blowup (as in $P=BPP$), it is far from clear that this blowup can be made $n^{o(1)}$ (see [15] for the state-of-the-art on such results). Can we provide a positive answer to Question 1 with respect to deterministic algorithms by designing deterministic Direct-WTERs?
Additional motivation comes from the hope of using the expander decomposition method in order to get breakthrough derandomization results, along the lines of Question 2, for problems where the current randomized algorithms are much faster than the current deterministic algorithms. Indeed, deterministic expander decompositions [16] have already played a major role in some of the most remarkable derandomization results of recent years, e.g. for Global Min-Cut [29, 37, 31, 30, 25].
@@
The following problems admit deterministic Direct-WTERs: Maximum Matching, Minimum Vertex Cover, $k$-Clique Detection, $k$-Clique Counting, Max-Clique, Max-Cut, Minimum dominating Set, and $H$-Subgraph Detection ($m=\tilde{O}(m)$ and $H$ does not contain pendant vertices).
@@
An important feature of our deterministic core gadget is that it remains simple and efficient. Interestingly, this stands in contrast with other derandomization results in fine-grained complexity, which often tends to involve sophisticated methods and some slowdown (see, e.g.,  [11, 18]). As discussed earlier, simplicity is an important aspect of Direct-WTERs, not just because it makes them more accessible to the community, but also because it strengthens the message that expander decompositions become useless in the presence of Direct-WTERs. In Section 2, we provide an intuitive overview of how our new results are obtained by a modification to the core gadget of Abboud and Wallheimer [3]. Section 4 provides the construction itself and the full analysis.
Interestingly, our Direct-WTERs also improve upon the blowup in the number of added edges over the Direct-WTERs in [3], resulting in $\Omega(1)$-expanders with $O(n)$ vertices and $O(m+n)$ edges, whereas in [3], the expanders have $O(m+n\log n)$ edges. Note that this blowup is optimal since any expander is connected and, therefore, must contain $\Omega(n)$ edges. We also demonstrate that our core gadget preserves the following graph properties: (1) Bipartiteness-preserving; we can modify the core gadget so that if $G$ is bipartite, then so is the expander, and (2) Degree-preserving; if the maximum degree in $G$ is $\Delta$, then the maximum-degree in the expander is $2\Delta+O(1)$.
@@
    For the exponential-time problems in Theorem 2: Minimum Vertex Cover, Minimum Dominating Set, Max-Clique, and exact Max-Cut, the blowup in the number of vertices in the output graph must be subject to stronger restrictions than for polynomial-time problems. In particular, the blowup should be $n+o(n)$, to show that expanders are worst-case instances.333 Otherwise, an exponential speed-up on expanders does not necessarily translate to an exponential speed-up on general graphs. For such problems, we employ a generalized core gadget, which gives a tradeoff between the conductance and the blowup, resulting in Direct-WTERs providing, for every $0<\varepsilon\leq 1$, conductance $\Omega(\varepsilon)$ and blowup $\varepsilon n$.
@@
    For the Max-Cut problem, there is an interesting related work on the approximation variant of the problem on expanders. A famous algorithm by Goemans and Williamson [20] obtains a $>0.878$-approximation for the maximum cut in general graphs, based on a Semidefinite Programming relaxation. In search of other, perhaps simpler methods for approximating the max-cut beyond the trivially obtained $1/2$-approximation [^4] Which follows from the fact that the maximum cut is at least $m/2$ in any graph with $m$ edges., Trevisan [42] posed the following question. Is there a combinatorial algorithm that achieves better than $1/2$-approximation?
    A positive answer to this question was given by Kale and Seshadhri [27], which remains the current-best combinatorial algorithm for this problem. A recent paper by Peng and Yoshida [36] addresses this question on expanders, providing a combinatorial algorithm for approximating the maximum cut on $\phi$-expanders. Namely, the authors provide an algorithm that given $\varepsilon$, computes a $(1/2+\varepsilon)$-approximation, subject to $\varepsilon=O(\phi^{2})$. Specifically, it computes a value $x$ such that $(1/2+\varepsilon)MC(G)\leq x\leq MC(G)$, where $MC(\cdot)$ denotes the cardinality of the maximum cut. Moreover, its running time is sublinear when $\phi$ is a constant. Our Direct-WTER for Max-Cut in Theorem 2, on the other hand, is a reduction that given $G$ and $0<\phi\leq 1$, outputs an $\Omega(\phi)$-expander $G_{exp}$, and the maximum cut in $G_{exp}$ is $MC(G_{exp})\leq(1+4\phi)MC(G)$, assuming the input graph is not too sparse [^5] More precisely, this holds if in $G$ there are $m>Cn$ edges for some constant $C$..
    Can we apply the Direct-WTER and then use the algorithm of Peng and Yoshida to get a combinatorial algorithm that $(1/2+\varepsilon)$-approximates the maximum cut in general graphs? Perhaps we may even improve upon the algorithm by Kale and Seshadhri, as both the Direct-WTER and the algorithm of Peng and Yoshida are very efficient. However, the approximate value we get from this approach is $(1/2+\varepsilon)MC(G)\leq x\leq(1+4\phi)MC(G)$, or equivalently $(1/2+\varepsilon)/(1+4\phi)MC(G)\leq x\leq MC(G)$. For this approximation ratio to be larger than $1/2$, we need to pick $\varepsilon>2\phi$, but recall the constraint $\varepsilon\leq O(\phi^{2})$. Hence, this approach fails. This is not so surprising, since our Direct-WTER is quite elementary so we do not expect to make use of it as a subroutine inside another algorithm. Instead, we can interpret this result as a limitation to algorithms for Max-Cut on $\phi$-expanders; that one cannot obtain a $(1/2+\varepsilon)$-approximation for some $22\phi<\varepsilon\leq 1/$ 2 unless this highly unlikely approach works.
@@ 1.2 Direct-WTERs in the Fully Dynamic setting
An important related work is a recent paper by Henzinger, Paz, and Sricharan [26] (abbreviated as HPS), who initiated the study of Question 1 in the dynamic model, regarding the complexity of fundamental problems on dynamic expanders. A dynamic expander is a dynamic graph that undergoes edge updates but remains an $\Omega(1)$-expander at any point in time. They adapt lower bound proofs from fine-grained complexity, so that they hold even on dynamic expanders. Their techniques differ from the self-reduction approach of Direct-WTERs and correspond to the first of the three methods outlined above to answer Question 1. In particular, they base their results on the Online Matrix Vector (OMv) conjecture, which was introduced by Henzinger et al. [24] to prove lower bounds for various dynamic problems. HPS obtained their results by adapting these lower bound proofs to the case of constant-degree expanders, proving that dynamic expanders whose maximum degree remains bounded by a constant are OMv-hard. The problems they consider in this context are Maximum Matching, Densest Subgraph, and $st$-Shortest Path (abbreviated as $st$-SP). This leaves us wondering with the following questions:
@@
    For the Densest Subgraph problem, Henzinger et al. [24] prove a lower bound of $n^{1/3-o(1)}$ per update under OMv for general graphs. By modifying their reduction, HPS were able to show a weaker $n^{1/4-o(1)}$ lower bound for expanders.777 We remark that the proof in Henzinger et al. [24, Corollary 3.26] gives an $n^{1/3-o(1)}$ lower bound under OMv, while the introductions of both [24] and HPS [26] mention, erroneously, an $n^{1/2-o(1)}$ lower bound.
    As a consequence of our direct DD-WTER, we conclude that the $n^{1/3-o(1)}$ lower bound for general graphs also holds for expanders. One subtlety towards this result is that the reduction of Henzinger et al. produces very sparse graphs with $m\leq 2n$ edges while our DD-WTER assumes that $m>42n$. In Appendix A, we discuss how to modify the original reduction so that denser graphs are produced. Another strength of the DD-WTER compared to HPS is that it is a self-reduction, hence it does not depend on the OMv Conjecture to get a lower bound (at least when the $m>42n$ assumption is made) [^8] We remark that the limitation of $m>42n$ in this approach is due to the fact that to augment a graph to become an expander requires at least adding some amount of edges, which unavoidably affects the densest subgraph..
@@
    For the Maximum Matching problem on graphs of maximum degree $O(n^{t})$, for any $0\leq t\leq 1$, we show that this problem is OMv-hard on $\Omega(1)$-expanders as well. This improves upon HPS, who only prove it for the constant-degree case (i.e. $t=0$). Our result is obtained by combining a lower bound proof from HPS for graphs of maximum degree $O(n^{t})$, with our degree-preserving DD-WTER for Maximum Matching.
@@
    We demonstrate that our core gadget is useful even outside the context of self-reductions by showing a DD-WTER for Graphical OMv, an equivalent graph formulation of the OMv problem. This implies that we can lift OMv-based lower bounds to $\Omega(1)$-expanders for many problems, in particular, for problems such as Maximum Matching and $st$-SP, which HPS considered. In particular, we demonstrate the power of this technique by proving that $st$-SP is OMv-hard on $\Omega(1)$-expanders. [^9] Note that in comparison with HPS, we do not prove hardness for graphs of constant degree, which is outside the scope of our paper.
@@
Finally, let us discuss another aspect of how our work and HPS differ. One of the main motivations for our work is to address Questions 2 and 3 about the applicability of expander decompositions in algorithms, whereas HPS main motivation is to gain a better understanding of the complexity of dynamic problems on various graph families, including expanders, along the lines of Question 1. As we have seen, DD-WTERs imply that expander decompositions are useless because they show equivalence between the worst-case and the expander-case complexities. We remark that the results of HPS imply the same, assuming their obtained lower bounds are tight. Hence, it implies a conditional answer to these questions, whereas self-reductions imply an unconditional one.
@@
The above results demonstrate that the impact of Direct-WTERs goes beyond the classic Word-RAM model of computation. A natural continuation of this is to apply these techniques to more models of computation, where expanders, expander decompositions, and derandomization are important subjects. In Section 8, we discuss the applicability of Direct-WTERs in distributed models of computation. We show that while there are limitations in adapting Direct-WTERs to CONGEST, it is possible to do so in CONGESTED-CLIQUE and MPC.

### v2

@@ Abstract
A recent paper by Abboud and Wallheimer [ITCS 2023] presents self-reductions for various fundamental graph problems, which transform worst-case instances to expanders, thus proving that the complexity remains unchanged if the input is assumed to be an expander. An interesting corollary of their self-reductions is that if some problem admits such reduction, then the popular algorithmic paradigm based on expander-decompositions is useless against it. In this paper, we improve their core gadget, which augments a graph to make it an expander while retaining its important structure. Our new core construction has the benefit of being simple to analyze and generalize while obtaining the following results:
@@
  • An extension of the results to other models of computation, such as the Fully Dynamic model and the Congested Clique model. In the former, we either improve or provide an alternative approach to some recent hardness results for dynamic expander graphs by Henzinger, Paz, and Sricharan [ESA 2022].
@@ 1 Introduction
It is possible to cook up problems for which the answer to 1 is negative. For instance, we can solve connectivity in constant time if the input is promised to be an expander, but it requires linear time in the worst case. A less obvious example with such gaps is counting spanning trees [31]. However, for many (perhaps most) interesting graph problems, the answer seems to be positive: expanders do not make the problem any easier. In other words, the expander-case is also worst-case. But how do we prove that? Let us discuss three methods and their drawbacks.
@@
    The main drawback with this approach is that we are interested in answering 1 even when (or rather, especially when) we have not already resolved the worst-case time complexity of our problem, in which case we do not even have a matching lower bound (e.g., the Maximum Matching problem). Another drawback is that when asking for the fine-grained complexity of problems, the existing lower bounds are usually conditioned on strong assumptions, and one may hope to get an unconditional answer to 1.
@@
    The second approach uses the expander decomposition method. This is a popular paradigm in recent years that suggests we can solve graph problems by (1) decomposing the graph into vertex-disjoint expanders with a small number of edges between them, (2) solving the problem on each expander separately, and (3) combining all the answers efficiently. Step (3) requires problem-specific techniques. If we can solve steps (1) and (3) for a problem, then we have effectively shown that any improvement on expanders will yield an improvement on worst-case graphs, giving a positive answer to 1. Efficient algorithms for computing such expander decompositions (for step (1)) are known both in the Word-RAM  [5, 37, 27, 39, 33] and in other models of computation including dynamic [37], distributed [12], and recently even in streaming [16]. Applications of this paradigm have led to many breakthroughs in recent years to problems such as Maximum Flow [13], Dynamic Connectivity [22], Gomory-Hu Trees [2], Minimum Spanning Trees [32], and Triangle Enumeration [12].
@@
    • The blowup in the number of vertices and edges in $G_{exp}$ is $|V_{exp}|\leq K$ and $|E_{exp}|\leq M$ for some $K:=K(n,m)$ and $M:=M(n,m)$.
@@
    The main contribution of Abboud and Wallheimer was to show that some fundamental problems, such as $k$-Clique Detection and Maximum Matching, admit simple Direct-WTERs. In particular, their Direct-WTERs make a graph an expander by employing a core gadget that augments it with $O(n)$ vertices and $O(m+n\log n)$ random edges and then applies additional gadgets that control the solution. In particular, they obtain a near-linear blowup. Their results are surprising because such Direct-WTERs do not employ any of the heavy machinery that usually comes with expander decompositions, yet they output quantitatively better expanders: the outputs of Direct-WTERs are $\Omega(1)$-expanders (by definition), whereas expander decompositions can only produce $O(1/\log n)$-expanders (that are not as expanding) [37, 6].
@@
Motivated by the appeal of the method of Direct-WTERs towards answering 1 and 2, our goal is to develop this theory further. Toward that, we address the two main limitations that were highlighted in [3]: (1) the Direct-WTERs are randomized whereas ED-WTERs are deterministic [15], and (2) the results are restricted to the Word-RAM model, whereas expander decompositions are popular tools in other models as well. In addition, we continue their line of work by providing Direct-WTERs to additional problems. Let us motivate these two topics before stating our results formally.
@@ 1.1 Deterministic Direct-WTERs
The randomized Direct-WTERs of Abboud and Wallheimer [3] prove that $\Omega(1)$-expanders are worst-case instances of many problems if we allow algorithms to be randomized. They leave us wondering if perhaps $\Omega(1)$-expanders are truly easier for deterministic algorithms. We remark that while it is believed that all algorithms can be derandomized by incurring a small polynomial blowup (as in $P=BPP$), it is far from clear that this blowup can be made $n^{o(1)}$ (see [14] for the state-of-the-art on such results). Can we provide a positive answer to 1 with respect to deterministic algorithms by designing deterministic Direct-WTERs?
Additional motivation comes from the hope of using the expander decomposition method in order to get breakthrough derandomization results, along the lines of 2, for problems where the current randomized algorithms are much faster than the current deterministic algorithms. Indeed, deterministic expander decompositions [15] have already played a major role in some of the most remarkable derandomization results of recent years, e.g. for Global Min-Cut [28, 36, 30, 29, 24].
@@
The following problems admit deterministic Direct-WTERs: Maximum Matching, Minimum Vertex Cover, $k$-Clique Detection, $k$-Clique Counting, Max-Clique, Max-Cut, Minimum Dominating Set, and $H$-Subgraph Detection ($m=\tilde{O}(n)$ and $H$ does not contain pendant vertices).
@@
An important feature of our deterministic core gadget is that it remains simple and efficient. Interestingly, this stands in contrast with other derandomization results in fine-grained complexity, which often tends to involve sophisticated methods and some slowdown (see, e.g.,  [10, 17]). As discussed earlier, simplicity is an important aspect of Direct-WTERs, not just because it makes them more accessible to the community, but also because it strengthens the message that expander decompositions become useless in the presence of Direct-WTERs. In Section 2, we provide an intuitive overview of how our derandomization is obtained by a modification to the core gadget of Abboud and Wallheimer [3] and in Section 4 we provide the construction itself.
Interestingly, our Direct-WTERs also improve upon the blowup in the number of added edges over the Direct-WTERs in [3], resulting in $\Omega(1)$-expanders with $O(n)$ vertices and $O(m+n)$ edges, whereas in [3], the expanders have $O(m+n\log n)$ edges. Note that this blowup is optimal since any expander is connected and, therefore, must contain $\Omega(n)$ edges. We also demonstrate that our core gadget preserves the following graph properties: (1) Bipartiteness-preserving; we can modify the core gadget so that if $G$ is bipartite, then so is the expander, and (2) Degree-preserving; if the maximum degree in $G$ is $\Delta$, then the maximum-degree in $G_{exp}$ is $2\Delta+O(1)$.
@@
    For the exponential-time problems in Theorem 2: Minimum Vertex Cover, Minimum Dominating Set, Max-Clique, and Max-Cut, the blowup in the number of vertices in the output graph must be subject to stronger restrictions than for polynomial-time problems. In particular, the blowup should be $n+o(n)$, to show that expanders are worst-case instances.222 Otherwise, an exponential speed-up on expanders does not necessarily translate to an exponential speed-up on general graphs. For such problems, we employ a generalized core gadget, which gives a tradeoff between the conductance and the blowup, resulting in Direct-WTERs providing, for every $0<\varepsilon\leq 1$, conductance $\Omega(\varepsilon)$ and blowup $\varepsilon n$.
@@
    For the Max-Cut problem, there is an interesting related work on the approximation variant of the problem on expanders. A famous algorithm by Goemans and Williamson [19] obtains a $>0.878$-approximation for the maximum cut in general graphs, based on a Semidefinite Programming relaxation. In search of other, perhaps simpler methods for approximating the max-cut beyond the trivially obtained $1/2$-approximation [^3] Which follows from the fact that the maximum cut is at least $m/2$ in any graph with $m$ edges., Trevisan [40] posed the following question: Is there a combinatorial algorithm that achieves better than $1/2$-approximation?
    A positive answer to this question was given by Kale and Seshadhri [26], which remains the current-best combinatorial algorithm for this problem. A recent paper by Peng and Yoshida [35] addresses this question on expanders, providing a combinatorial algorithm for approximating the maximum cut on $\phi$-expanders. Namely, the authors provide an algorithm that given $\varepsilon$, computes a $(1/2+\varepsilon)$-approximation, subject to $\varepsilon=O(\phi^{2})$. In more detail, it computes a value $x$ such that $(1/2+\varepsilon)MC(G)\leq x\leq MC(G)$, where $MC(\cdot)$ denotes the cardinality of the maximum cut in $G$. Moreover, its running time is sublinear when $\phi$ is a constant. Our Direct-WTER for Max-Cut in Theorem 2, on the other hand, is a reduction that given $G$ and $0<\phi\leq 1$, outputs an $\Omega(\phi)$-expander $G_{exp}$, and the maximum cut in $G_{exp}$ is $MC(G_{exp})\leq(1+4\phi)MC(G)$, assuming the input graph is not too sparse, say, $m=\omega(n)$.
    Can we apply the Direct-WTER and then use the algorithm of Peng and Yoshida to get a combinatorial algorithm that $(1/2+\varepsilon)$-approximates the maximum cut in general graphs? Perhaps we may even improve upon the algorithm by Kale and Seshadhri, as both the Direct-WTER and the algorithm of Peng and Yoshida are very efficient. However, the approximate value we get from this approach is $(1/2+\varepsilon)MC(G)\leq x\leq(1+4\phi)MC(G)$, or equivalently $(1/2+\varepsilon)/(1+4\phi)MC(G)\leq x\leq MC(G)$. For this approximation ratio to be larger than $1/2$, we need to pick $\varepsilon>2\phi$, but recall the constraint $\varepsilon\leq O(\phi^{2})$. Hence, this approach fails. This is not so surprising, since our Direct-WTER is quite elementary so we do not expect to make use of it as a subroutine inside another algorithm. Instead, this result should be interpreted as a limitation to algorithms for Max-Cut on $\phi$-expanders; that one cannot obtain a $(1/2+\varepsilon)$-approximation for some $22\phi<\varepsilon\leq 1/$ 2 unless this (unlikely) approach works.
@@ 1.2 Direct-WTERs in the Fully Dynamic setting
An important related work is a recent paper by Henzinger, Paz, and Sricharan [25] (abbreviated as HPS), who initiated the study of 1 in the dynamic model, regarding the complexity of fundamental problems on dynamic expanders. A dynamic expander is a dynamic graph that undergoes edge updates but remains an $\Omega(1)$-expander at any point in time. They adapt lower bound proofs from fine-grained complexity, so that they hold even on dynamic expanders. Their techniques differ from the self-reduction approach of Direct-WTERs and correspond to the first of the three methods outlined above to answer 1. In particular, they base their results on the Online Matrix Vector (OMv) conjecture, which was introduced by Henzinger et al. [23] to prove lower bounds for various dynamic problems. HPS obtained their results by adapting these lower bound proofs to the case of constant-degree expanders, proving that dynamic expanders whose maximum degree remains bounded by a constant are OMv-hard. The problems they consider in this context are Maximum Matching, Densest Subgraph, and $st$-Shortest Path (abbreviated as $st$-SP). This leaves us wondering with the following questions:
@@
    For the Densest Subgraph problem, Henzinger et al. [23] prove a lower bound of $n^{1/3-o(1)}$ per update under OMv for general graphs. By modifying their reduction, HPS were able to show a weaker $n^{1/4-o(1)}$ lower bound for expanders.555 We remark that the proof in Henzinger et al. [23, Corollary 3.26] gives an $n^{1/3-o(1)}$ lower bound under OMv, while the introductions of both [23] and HPS [25] mention, erroneously, an $n^{1/2-o(1)}$ lower bound.
    As a consequence of our DD-WTER, we conclude that the $n^{1/3-o(1)}$ lower bound for general graphs also holds for expanders. One subtlety towards this result is that the reduction of Henzinger et al. produces very sparse graphs with $m\leq 2n$ edges while our DD-WTER assumes that $m>42n$. In Appendix A, we discuss how to modify the original reduction so that denser graphs are produced. Another strength of the DD-WTER compared to HPS is that it is a self-reduction, hence it does not depend on the OMv Conjecture to get a lower bound (at least when the $m>42n$ assumption is made) [^6] We remark that the limitation of $m>42n$ in this approach is due to the fact that to augment a graph to become an expander requires at least adding some amount of edges, which unavoidably affects the densest subgraph..
@@
    For the Maximum Matching problem on graphs of maximum degree $O(n^{t})$, for any $0\leq t\leq 1$, we show that this problem is OMv-hard on $\Omega(1)$-expanders as well. This improves upon HPS, who only prove it for the constant-degree case (i.e. $t=0$[25, Theorem 12], while implicitly leaving an open question regarding $t>0$. While intuitively, the constant-degree case should be the most difficult to prove a lower bound for, it is not immediate, as techniques that artificially increase the degrees in the graph (e.g. attaching a star to every vertex) tend to also increase the number of vertices, thus resulting in weaker lower bounds. Instead, our result is obtained by combining a lower bound by HPS for graphs (not expanders) of maximum degree $O(n^{t})$ [25, Theorem 12], with our degree-preserving DD-WTER for Maximum Matching. Hence, we get the following corollary:
@@
    We demonstrate that our core gadget is useful even outside the context of self-reductions by showing a DD-WTER for Graphical OMv, an equivalent graph formulation of the OMv problem. This implies that we can lift OMv-based lower bounds to $\Omega(1)$-expanders for many problems, in particular, for problems such as Maximum Matching and $st$-SP, which HPS considered. We demonstrate the power of this technique by proving that $st$-SP is OMv-hard on $\Omega(1)$-expanders. [^7] Note that in comparison with HPS, we do not prove hardness for graphs of constant degree, which is outside the scope of our paper.
@@
Finally, let us emphasize another aspect of how our work and HPS differ. One of the main motivations for our work is to address 2 and 3 about the applicability of expander decompositions in algorithms, whereas HPS main motivation is to gain a better understanding of the complexity of dynamic problems on various graph families, including expanders, along the lines of 1. As we have seen, DD-WTERs imply that expander decompositions are useless because they show equivalence between the worst-case and the expander-case complexities. We remark that the results of HPS imply the same, assuming their obtained lower bounds are tight. Hence, it implies a conditional answer to these questions, whereas self-reductions imply an unconditional one.
@@
The above results demonstrate that the impact of Direct-WTERs goes beyond the classic Word-RAM model of computation. A natural continuation of this is to apply these techniques to more models of computation, where expanders, expander decompositions, and derandomization are important subjects. In Section 8, we discuss the applicability of Direct-WTERs in distributed models of computation. We show that while there are limitations in adapting Direct-WTERs to CONGEST, it is possible to do so in CONGESTED-CLIQUE and MPC (Massively Parallel Computation).